# A Rose by Any Other Name: LLM-Generated Explanations Are Good Proxies for Human Explanations to Collect Label Distributions on NLI

### v1

@@ Abstract
Disagreement in human labeling is ubiquitous, and can be captured in human judgment distributions (HJDs). Recent research has shown that explanations provide valuable information for understanding human label variation (HLV) and large language models (LLMs) can approximate HJD from a few human-provided label-explanation pairs. However, collecting explanations for every label is still time-consuming. This paper examines whether LLMs can be used to replace humans in generating explanations for approximating HJD. Specifically, we use LLMs as annotators to generate model explanations for a few given human labels. We test ways to obtain and combine these label-explanations with the goal to approximate human judgment distribution. We further compare the resulting human with model-generated explanations, and test automatic and human explanation selection. Our experiments show that LLM explanations are promising for NLI: to estimate HJD, generated explanations yield comparable results to human’s when provided with human labels. Importantly, our results generalize from datasets with human explanations to i) datasets where they are not available and ii) challenging out-of-distribution test sets.
@@ 1 Introduction
Figure 1: Recent research has shown that LLMs can approximate human judgment distribution (HJD) in natural language inference (NLI) with the help of human explanations, as shown in the upper part. While human explanations are still relatively expensive and scarce in most datasets, we ask in the lower part: Can LLMs provide reasonable generated explanations for different NLI labels to approximate HJD?
@@
Recent studies have found that LLMs can effectively provide explanations for tasks such as reasoning, sentiment analysis, and even business processes Li et al. (2022); Huang et al. (2023); Fahland et al. (2024). We instead study automatic explanation generation for the NLI task. In this paper, we investigate if LLMs could provide reasonable explanations for NLI instance labels, and whether the generated explanations are of sufficient quality to replace costly human-provided explanations. Our key idea is shown in Figure 1. Specifically, we let LLMs generate model explanations supporting each NLI label, respectively. We first examine if LLMs with diverse model explanations but without multiple human labels are good enough in approximating HJD. The result is positive but does still fall short of the performance of LLMs with human explanations. We then consider using a few human labels to guide the selection of model explanations to help generate MJDs. MJDs from LLM and model explanations result in comparable scores with MJDs from LLM and human explanations —“A rose by any other name would smell as sweet.” [^1] A quote from Romeo and Juliet used to metaphorically argue the intrinsic qualities or nature of something remain the same, regardless of its name or origin. Furthermore, we extend this method to datasets without explanations, showing our method generalizes to this more common scenario.
@@
  • A human annotation study and ablation show that explanation variability may serve as a potential indicator for evaluating HLV, and the relevancy of explanations is crucial.
@@ 2 Generating Model Explanation
Collecting human explanations for an annotation decision is labor-intensive and missing in most Natural Language Inference (NLI) datasets. In this paper, we intend to examine whether Large Language Models (LLMs) can replace humans in generating these explanations. But how do we best query and select resulting model explanations? In this section, we detail the generation of model explanations, the selection of label-free and label-guided explanations, as illustrated in Figure 2.
@@
We prompt LLMs to generate explanations for specific premise and hypothesis pairs and a given NLI label (see Table 6 for details). Since NLI is subject to variation within the label Bowman et al. (2015); Jiang et al. (2023); Weber-Genzel et al. (2024) and multiple annotators can agree on the same label for different reasons, we ask LLMs to list all possible explanations for each NLI label for a given instance. We next introduce two explanation selection strategies: label-free and label-guided.
Figure 2: Illustration of the process of generating model explanation using longest explanations. The label-free scenario uniformly selects one explanation per {E, N, C} label, whereas the label-guided approach follows human NLI labels to select three longest E, one longest N, and one longest C.
@@
A few NLI datasets have addressed the issue of human label variation (HLV) by including a small number of label annotations on the same instances, 5 for MNLI Williams et al. (2018) and 4 for VariErr Weber-Genzel et al. (2024). Although earlier usages of these datasets are centered around the final label aggregated by majority voting, we consider using the small number of human labels as guidance to help build a combination of model explanations to approximate HJD. Unlike the label-free approach, we select explanations based on the annotated NLI labels for each instance. For example, three explanations for Entailment, one for Neutral, and one for Contradiction are selected in Figure 2.
@@
Since LLMs are prompted to exhaustively output explanations for given instances and labels, we propose two modes for selecting a desired number of explanations: one based on the linear order of LLM outputs (first) and another based on the length of the output explanations (longest). For example, if two annotators have annotated an NLI instance with Entailment, we select the two longest model explanations that support Entailment under the longest mode; or the initial two output Entailment explanations under the first mode. The first mode represents the primary preferences of LLMs, particularly in cases when prompting without explicit requirements to output all possible explanations. The longest mode can reveal more information regarding the reasoning between the premise and the hypothesis. [^3] Note that the overall overlap rate between first and longest explanations are about 18.9%. See Table 10 in Appendix for detailed statistics. Our experiments reveal that first and longest modes achieve similar results, cf. §6. Thus, in the main paper we report performances using longest explanation(s); results with the first mode are in the appendix.
@@ 3 Can Model Explanations Help LLMs Approximate HJD as Humans Do?
Our first research question (RQ1) concerns whether LLM-generated explanations can model human judgment distribution (HJD) as effectively as the human-written explanations.
@@ 3.1 Experimental Setup
ChaosNLI Nie et al. (2020b) includes 100 crowd-sourced annotations per instance and is considered gold Human Label Distribution (HJD) in our experiments. VariErr Weber-Genzel et al. (2024), which tackles the explainability of NLI by asking a few experts to record the explanation behind each NLI label explicitly, includes 4 label and explanation annotations per instance on 341 items that overlap with ChaosNLI and MNLI (Williams et al., 2018). The overlapping subset of MNLI Williams et al. (2018) includes 5 label annotations per instance but without human-written explanations. Details in Table 8.
@@ 3.2 Results
Figure 3: Llama3 Visualization in the ternary plot Gruber et al. (2024). Each point represents the label distribution for one NLI instance. From left to right listed ChaosNLI HJD (red), Llama3 (blue), Llama3 with human explanations (blue), Llama3 with Label-Free (LF), Varierr/MNLI Label-Guided (LG) model explanations (3$\times$×purple).
@@
For model explanations, we first want to explore whether eliminating human labeling and only entering one model explanation for each of the three possible classes is enough for LLMs to approximate the HJD. The results show that Label-Free explanations perform poorly in distance comparisons when compared to LLMs, most likely due to the equal distribution of three labels. However, the Label-Free method still scores significantly higher in the fine-tuning comparison and on D.Corr, and the latter two better reflect practical performance and global correlation. As the Label-Free performance is between LLMs and LLMs with human explanations (including VariErr labels), results show that model explanations do provide useful HLV information to allow LLMs to generate better MJDs. Figure 3d plots the label-free (LF) MJD of Llama3. Its distribution is relatively smooth, uniform, and less centered on particular labels than the original MJD of Llama3 in Figure 3b. We also visualized the MJDs from GPT-4o. Unlike Llama3, the original MJD of GPT-4o in Figure 4a is relatively uniform, and Label-Free model explanations help move some wrong points upward (in NLI, more disagreements are found between E-N and N-C, and fewer are born in E-C, the bottom points in the triangle). This may explain why Label-Free can be better than LLMs on the fine-tuning comparison and D.Corr.
@@
To investigate the generalizability of our approach, [^4] Considering the experimental cost, we only tested this setting on open-sourced LLMs. Mixtral results in Appendix B.3. we used MNLI-guided explanation generation utilizing the 5 NLI labels for each MNLI instance in the overlapping NLI data subset. [^5] A preliminary experiment on 59 NLI instances with 5 explanation-label pairs from VariErr found that varying number of explanations between 3 to 5 does not have much impact. Table 1 shows that MNLI-guided achieves the best MJD on F1 for both BERT and RoBERTa FT, as well as on D.Corr. Figure 3f visualizes the MJD of MNLI-guided, and its distribution is more similar to Chaos NLI HJD. Both are smoother at the upper corner of the triangle than human (Figure 3c) or VariErr-guided explanations (Figure 3e), and are also more skewed towards the contradiction side than Label-Free (Figure 3d) and Llama3 itself (Figure 3b). The comparable performance of VariErr and MNLI-guided explanations to human explanations shows the scalability of our model-generated explanation.
4 Can Model-Generated Explanations Enhance Performance on OOD Task?
@@ 4.1 Experiment Setup
The ANLI dataset  Nie et al. (2020a) is a challenging NLI dataset collected by an adversarial procedure. Mechanical Turkers are instructed to continue writing hypotheses for a given context and target label until a trained BERT/RoBERTa model (using MNLI, SNLI, etc.) outputs a wrong label prediction. This iterative process is conducted on three rounds (R1-R3) of annotations, and each round contains different context texts, mainly from Wikipedia, but R3 includes additional news, fiction, speech, and other contexts. We conduct OOD evaluation on R1-R3 data of the ANLI test set.
@@
Table 2: ANLI test results. Scores reported in the table are Weighted F1 scores. The bold numbers indicate the best results under the corresponding LLM. Detailed results for individual runs as well as Mixtral’s performance are elaborated in Appendix C.
@@ 5 Human versus Model Explanations: Are They Different and Does It Matter?
All comparisons above treat the ChaosNLI HJD as the comparison target to explore RQ3. It would be useful to temporarily alter our perspective and treat the MJD with human explanations as the target, highlighted in green in Table 3. We observe that on all metrics, Llama3 without explanation and with Label-Free are far from MJD with human explanations. Gradually replacing by model explanations keeps the generated MJDs slightly away from the centric but remains very similar. [^7] We examine the similarities between model/human explanations following Giulianelli et al. (2023) (Table 17). Our observation mirrors theirs in that model explanations differ moderately from human explanations regarding lexicon, syntax, and semantics. Nevertheless, LLMs still found a way to obtain comparable information for modeling HJD. It is worth noting that although VariErr’s label distribution is far considering traditional instance-level distance comparison metrics, it is relatively similar in the global metric D.Corr. This observation is consistent with that VariErr distribution performs well in FT comparison in Table 1, further corroborating the finding by Chen et al. (2024) on the suitability of D.Corr in comparing label distributions.
@@ 6 Can Human Preference Lead to Better Explanation Selection?
Results are in Table 4 (details in Appenidx E). Surprisingly, model explanations from the unpreferred set achieved the best results on most metrics, which is different from our original expectation. This may be due to the unpreferred explanations being more diverse than the preferred ones (note that this is limited to the judgment of a single annotator). To verify this hypothesis, we average similarities among each pair of explanations ($C\binom{2}{n}$ pairs for n explanations) on each NLI instance. Table 5 shows that human explanations have the greatest variability on all of Giulianelli et al. (2023) similarity measures (the lower the value, the higher the variability). Moreover, model explanations that counter the preferences of one human annotator (unpreferred) have a higher variability, providing more diverse perspectives, which aligns with the observation in Table 4. These experiments show the potential of variability as a metric for measuring the model explanations when helping LLMs approximate HJD. We expanded this variability check to the main experiment. Results in Table 20 show that MNLI-guided explanations from Llama3 have the best variability, which is consistent with the results in Table 1. However, variability cannot be directly linked to the main results in all circumstances, cf. Appendix E for a discussion.
@@ 7 Related Work
Recently LLM-generated explanations have been used in various tasks (e.g., reasoning, sentimental analysis, recommender systems, education, abusive language detection. Li et al., 2022; Huang et al., 2023; Lubos et al., 2024; Abu-Rasheed et al., 2024; Di Bonaventura et al., 2024). Kunz and Kuhlmann (2024) find that LLM-generated explanations show selectivity and contain illustrative elements, but less frequently are subjective or misleading. Unlike previous methods, we guide LLMs to generate more diverse explanations for NLI to analyze HLV.
@@ Acknowledgements
We thank the members of the MaiNLP lab for their insightful feedback on earlier drafts of this paper. We specifically appreciate the suggestions of Xinpeng Wang, Robert Litschko, Philipp Mondorf, Bolei Ma, Soh-Eun Shim, Felicia Körner, Andreas Säuberli, Verena Blaschke, Shijia Zhou, Diego Frassinelli, Yang Janet Liu, Michael A. Hedderich, and Rob van der Goot. We extend our gratitude to Huangyan Shan for annotating the human preference experiment. BC acknowledges his membership in the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program. We are also grateful to the anonymous reviewers for their constructive feedback. This research is supported by ERC Consolidator Grant DIALECT 101043235.

### v2

@@ Abstract
Disagreement in human labeling is ubiquitous, and can be captured in human judgment distributions (HJDs). Recent research has shown that explanations provide valuable information for understanding human label variation (HLV) and large language models (LLMs) can approximate HJD from a few human-provided label-explanation pairs. However, collecting explanations for every label is still time-consuming. This paper examines whether LLMs can be used to replace humans in generating explanations for approximating HJD. Specifically, we use LLMs as annotators to generate model explanations for a few given human labels. We test ways to obtain and combine these label-explanations with the goal to approximate human judgment distributions. We further compare the resulting human with model-generated explanations, and test automatic and human explanation selection. Our experiments show that LLM explanations are promising for NLI: to estimate HJDs, generated explanations yield comparable results to human’s when provided with human labels. Importantly, our results generalize from datasets with human explanations to i) datasets where they are not available and ii) challenging out-of-distribution test sets.
@@ 1 Introduction
Figure 1: Recent research has shown that LLMs can approximate a human judgment distribution (HJD) in natural language inference (NLI) with the help of human explanations, as shown in the upper part. While human explanations are still relatively expensive and scarce in most datasets, we ask in the lower part: Can LLMs provide reasonable generated explanations for different NLI labels to approximate HJD?
@@
Recent studies have found that LLMs can effectively provide explanations for tasks such as reasoning, sentiment analysis, and even business processes Li et al. (2022); Huang et al. (2023); Fahland et al. (2024). We instead study automatic explanation generation for the NLI task. In this paper, we investigate if LLMs could provide reasonable explanations for NLI instance labels, and whether the generated explanations are of sufficient quality to replace costly human-provided explanations. Our key idea is shown in Figure 1.
Specifically, we let LLMs generate model explanations supporting each NLI label, respectively. We first examine ($\mathrm{i}$) can LLMs approximate HJD via diverse model explanations without access to label annotations? The result is positive but does still fall short of the performance of LLMs with human explanations. We then consider using a few human labels to guide the selection of model explanations to help generate MJDs, and ask ($\mathrm{ii}$) are model explanations comparable to human’s when helping LLMs to approximate HJD? Experiments show that MJDs from LLMs and model explanations result in comparable scores with MJDs from LLM and human explanations —“A rose by any other name would smell as sweet.” [^1] A quote from Romeo and Juliet used to metaphorically argue the intrinsic qualities or nature of something remain the same, regardless of its name or origin. Furthermore, we ask ($\mathrm{iii}$) does our approach extend to NLI datasets that do not have human-provided explanations? We extend the proposed method to datasets without explanations, showing our method generalizes to this more common scenario.
@@
  • A human annotation study and ablation shows that explanation variability may serve as a potential indicator for evaluating HLV, and the relevancy of explanations is crucial.
@@ 2 Generating Model Explanation
Collecting human explanations for an annotation decision is labor-intensive and missing in most Natural Language Inference (NLI) datasets. In this paper, we intend to examine whether Large Language Models can replace humans in generating these explanations. But how do we best query and select resulting model explanations? In this section, we detail the generation of model explanations, the selection of label-free and label-guided explanations, as illustrated in Figure 2.
@@
We prompt LLMs to generate explanations for specific premise and hypothesis pairs and a given NLI label (see Table 6 in Appendix A for details). Since NLI is subject to variation within the label Bowman et al. (2015); Jiang et al. (2023); Weber-Genzel et al. (2024) and multiple annotators can agree on the same label for different reasons, we ask LLMs to list all possible explanations for each NLI label for a given instance. We next introduce two explanation selection strategies: label-free and label-guided.
Figure 2: Illustration of the process of generating model explanation using the longest explanation selection heuristic. The label-free scenario uniformly selects one explanation per {E, N, C} label, whereas the label-guided approach follows human NLI labels to select three longest E, one longest N, and one longest C.
@@
A few NLI datasets have addressed the issue of human label variation (HLV) by including a small number of label annotations on the same instances, 5 for MNLI Williams et al. (2018) and 4 for VariErr Weber-Genzel et al. (2024). Although earlier usages of these datasets are centered around the final label aggregated by majority voting, we consider using the small number of human labels as guidance to help build a combination of model explanations to approximate HJDs. Unlike the label-free approach, we select explanations based on the annotated NLI labels for each instance. For example, for a given instance three explanations for Entailment, one for Neutral, and one for Contradiction are selected in Figure 2.
@@
Since LLMs are prompted to exhaustively output explanations for given instances and labels, we propose two modes for selecting a desired number of explanations: one based on the linear order of LLM outputs (first) and another based on the length in tokens of the output explanations (longest). For example, if two annotators have annotated an NLI instance with Entailment, we select the two longest model explanations that support Entailment under the longest mode; or the initial two output Entailment explanations under the first mode. The first mode represents the primary preferences of LLMs, particularly in cases when prompting without explicit requirements to output all possible explanations. The longest mode can reveal more information regarding the reasoning between the premise and the hypothesis. [^3] Note that the overall overlap rate between first and longest explanations are about 18.9%. See Table 10 in Appendix B.2 for detailed statistics. Our experiments reveal that first and longest modes achieve similar results, cf. §6. Thus, in the main paper we report performances using longest explanation(s); results with the first mode are in the appendix.
@@ 3 Can Model Explanations Help LLMs Approximate HJD as Humans Do?
Our first research question (RQ1) concerns whether LLM-generated explanations can model a human judgment distribution (HJD) as effectively as the human-written explanations.
@@ 3.1 Experimental Setup
ChaosNLI Nie et al. (2020b) includes 100 crowd-sourced annotations per instance and is considered gold Human Label Distribution (HJD) in our experiments. VariErr Weber-Genzel et al. (2024), which tackles the explainability of NLI by asking a few experts to record the explanation behind each NLI label explicitly, includes 4 label and explanation annotations per instance on 341 items that overlap with ChaosNLI and MNLI (Williams et al., 2018). The overlapping subset of MNLI Williams et al. (2018) includes 5 label annotations per instance but without human-written explanations. Details are provided in The Appendix B.1 in Table 8.
@@ 3.2 Results
Figure 3: Llama3 Visualization using a ternary plot Gruber et al. (2024). Each point represents the label distribution for one NLI instance. From left to right: ChaosNLI HJD (red), Llama3 (blue), Llama3 with human explanations (blue), Llama3 with Label-Free (LF), Varierr/MNLI Label-Guided (LG) model explanations (3$\times$×purple).
@@
For model explanations, we first explore whether eliminating human labeling and only entering one model explanation for each of the three possible classes is enough for LLMs to approximate the HJD. The results show that Label-Free explanations perform poorly in distance comparisons when compared to LLMs, most likely due to the equal distribution of the three labels. However, the Label-Free method still scores significantly higher in the fine-tuning comparison and on D.Corr, and the latter two better reflect practical performance and global correlation. As the Label-Free performance is between LLMs and LLMs with human explanations (including VariErr labels), results show that model explanations do provide useful HLV information to allow LLMs to generate better MJDs. Figure 3d plots the Label-Free (LF) MJD of Llama3. Its distribution is relatively smooth, uniform, and less centered on particular labels than the original MJD of Llama3 in Figure 3b. We also visualized the MJDs from GPT-4o. Unlike Llama3, the original MJD of GPT-4o in Figure 4a is relatively uniform, and Label-Free model explanations help move some wrong points upward (in NLI, more disagreements are found between E-N and N-C, and fewer are found between E-C, the bottom points in the triangle). This may explain why Label-Free can be better than LLMs on the fine-tuning comparison and D.Corr.
@@
To investigate the generalizability of our approach, [^4] Considering the experimental cost, we only tested this setting on open-sourced LLMs. Mixtral results in Appendix B.3. we used MNLI-guided explanation generation utilizing the 5 NLI labels for each MNLI instance in the overlapping NLI data subset. [^5] A preliminary experiment on 59 NLI instances with 5 explanation-label pairs from VariErr found that varying number of explanations between 3 to 5 does not have much impact. Table 1 shows that MNLI-guided achieves the best MJD on F1 for both BERT and RoBERTa FT, as well as on D.Corr. Figure 3f visualizes the MJD of MNLI-guided, and its distribution is more similar to Chaos NLI HJD. Both are smoother at the upper corner of the triangle than human (Figure 3c) or VariErr-guided explanations (Figure 3e), and are also more skewed towards the contradiction side than Label-Free (Figure 3d) and Llama3 itself (Figure 3b). The comparable performance of VariErr and MNLI-guided explanations to human explanations shows the scalability of our model-generated explanations.
4 Can Model-Generated Explanations Enhance Performance on OOD Data?
@@ 4.1 Experiment Setup
The ANLI dataset  Nie et al. (2020a) is a challenging NLI dataset collected by an adversarial procedure. Mechanical Turkers are instructed to continue writing hypotheses for a given context and target label until a trained BERT/RoBERTa model (using MNLI, SNLI, etc.) outputs a wrong label prediction. This iterative process is conducted on three rounds (R1-R3) of annotations, and each round contains different context texts, mainly from Wikipedia, but R3 includes additional news, fiction, speech, and other contexts. We conduct OOD evaluation on R1-R3 data of the ANLI test set. [^6] We discuss the OOD definition and data contamination risk of ANLI in Appendix F.
@@
Table 2: ANLI test results as Weighted F1 scores. Bold numbers indicate the best results under the corresponding LLM. Detailed results for individual runs as well as Mixtral’s performance are elaborated in Appendix C.
@@ 5 Human versus Model Explanations: Are They Different and Does It Matter?
All comparisons above treat the ChaosNLI HJD as the comparison target to explore RQ3. It would be useful to temporarily alter our perspective and treat the MJD with human explanations as the target, highlighted in green in Table 3. We observe that on all metrics, Llama3 without explanation and with Label-Free are far from MJD with human explanations. Gradually replacing by model explanations keeps the generated MJDs slightly away from the centric but remains very similar. [^8] We examine the similarities between model/human explanations following Giulianelli et al. (2023) (Table 17 in Appendix D). Our observation mirrors theirs in that model explanations differ moderately from human explanations regarding lexicon, syntax, and semantics. Nevertheless, LLMs still found a way to obtain comparable information for modeling HJD. It is worth noting that although VariErr’s label distribution is far considering traditional instance-level distance comparison metrics, it is relatively similar in the global metric D.Corr. This observation is consistent with that VariErr distribution performs well in FT comparison in Table 1, further corroborating the finding by Chen et al. (2024) on the suitability of D.Corr in comparing label distributions.
@@ 6 Can Human Preference Lead to Better Explanation Selection?
Results are in Table 4 (details in Appendix E). Surprisingly, model explanations from the unpreferred set achieved the best results on most metrics, which is different from our original expectation. This may be due to the unpreferred explanations being more diverse than the preferred ones (note that this is limited to the judgment of a single annotator). To verify this hypothesis, we average similarities among each pair of explanations ($C\binom{2}{n}$ pairs for $n$ explanations) on each NLI instance. Table 5 shows that human explanations have the greatest variability on all using Giulianelli et al. (2023)’s similarity measures (the lower the value, the higher the variability). Moreover, model explanations that counter the preferences of one human annotator (unpreferred) have a higher variability, providing more diverse perspectives, which aligns with the observation in Table 4. These experiments show the potential of variability as a metric for measuring the model explanations when helping LLMs approximate HJD. We expanded this variability check to the main experiment. Results in Table 20 in Appendix E show that MNLI-guided explanations from Llama3 have the best variability, which is consistent with the results in Table 1. However, variability cannot be directly linked to the main results in all circumstances, cf. Appendix E for a discussion.
@@ 7 Related Work
Recently LLM-generated explanations have been used in various tasks (e.g., reasoning, sentimental analysis, recommender systems, education, abusive language detection; Li et al., 2022; Huang et al., 2023; Lubos et al., 2024; Abu-Rasheed et al., 2024; Di Bonaventura et al., 2024). Kunz and Kuhlmann (2024) find that LLM-generated explanations show selectivity and contain illustrative elements, but less frequently are subjective or misleading. Wang et al. (2024) utilize human feedback on LLM-generated explanations to assist in collaborative annotation. This is even a potential future direction for us—to help build a human-LLM collaboration explanation NLI dataset. Wiegreffe et al. (2022) and Mishra et al. (2024) focus on evaluating the reasonableness, conciseness, and other aspects of LLM-generated explanations from a human-label-guided assessment perspective. Unlike previous methods, we guide LLMs to generate more diverse explanations for NLI to analyze HLV.
@@ 8 Conclusion
Besides, we also discuss the limitation of the explanation selection strategy presented in §2. The "first" explanation was chosen because it could reflect the initial intuition or immediate reasoning that a model or human might provide, offering a baseline for interpretability. The "longest" explanation was included to capture more detailed or comprehensive reasoning, which may provide deeper insights into the model’s decision-making process. While these characteristics serve as a useful starting point for analysis, we acknowledge that they represent only a subset of the dimensions that could be explored when evaluating explanations. Thus, how explanation diversity and quality influence the performance of HJD approximation is still up to future research.
@@ Acknowledgements
We thank the members of the MaiNLP lab for their insightful feedback on earlier drafts of this paper. We specifically appreciate the suggestions of Xinpeng Wang, Robert Litschko, Philipp Mondorf, Bolei Ma, Soh-Eun Shim, Felicia Körner, Andreas Säuberli, Verena Blaschke, Shijia Zhou, Diego Frassinelli, Yang Janet Liu, Michael A. Hedderich, and Rob van der Goot. We extend our gratitude to Huangyan Shan for annotating the human preference experiment. We are also grateful to the anonymous reviewers for their constructive feedback. BC acknowledges his membership in the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program. This research is supported by ERC Consolidator Grant DIALECT 101043235 and the UK Research and Innovation (UKRI) Frontier Research Grant EP/Y031350/1 EQUATE (the UK government’s funding guarantee for ERC Advanced Grants).